# SplatCo: Structure-View Collaborative Gaussian Splatting for Detail-Preserving Rendering of Large-Scale Unbounded Scenes

### v3

@@ V-E2 Effectiveness of Cross-View Assisted Training Strategy
Figure 8: Qualitative comparison of ablation study on cross-view assisted training strategy. Red & green squares highlight the visual differences for clearer comparison. By reconstructing the fine-grained structures in two scenes, such as the steel frame and signboards in the first row of images, and lane markings and street lamps in the second row of images, we demonstrate the superiority of the comprehensive surface completion training strategy and validate the effectiveness of its individual components.Please zoom in to view detailed results.
V-E3 Analysis of Hyperparameter Setting
In this section, we conduct a hyperparameter sensitivity analysis, focusing particularly on examining the impact of parameter $\tau$ on model performance. As shown in Table. VII, we evaluate five different configurations: $\tau\in\{0.4,0.6,0.8,1.0,1.2\}$. The experimental results demonstrate that the model achieves optimal performance when $\tau$ = 1.0. This analysis confirms the significant influence of parameter $\tau$ on the overall method performance.
TABLE VII: Ablation study on the setting of $\tau$.
method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
$\tau$=0.4 0.790 25.81 0.275
$\tau$=0.6 0.792 25.81 0.273
$\tau$=0.8 0.800 25.93 0.269
$\tau$=1.0 0.803 25.99 0.266
$\tau$=1.2 0.802 25.88 0.271
VI Conclusion and future work
In this paper, we introduce SplatCo, a structure–view collaborative Gaussian splatting framework designed for high-fidelity, large-scale unbounded scene rendering. We present a cross-structure collaboration module that synergistically fuses global tri-plane context with local grid features via hierarchical compensation, overcoming detail loss inherent in simpler representations. Then, we propose a cross-view assisted training strategy using gradient consensus and visibility-aware densification to address single-view overfitting and under-reconstruction issues prevalent in sparse large-scale captures, which stems from optimizing Gaussians based on individual viewpoints without explicit cross-view constraints. Furthermore, we enhance geometric integrity and generalization across views through a structural-consistency-guided pruning mechanism. Extensive experiments on 13 diverse large-scale scenes demonstrate the superiority of our approach in rendering quality and detail preservation compared to state-of-the-art methods.
Limitations: While the SplatCo method substantially improves detail and consistency in 3D reconstructions, its performance and practical application are subject to certain limitations. These limitations, however, also highlight critical avenues for future methodological enhancements.
Firstly, the efficacy of SplatCo remains sensitive to the input data characteristics. The initial quality of camera poses is a significant factor, as substantial inaccuracies can hinder optimal convergence. Furthermore, the density and overlap of viewpoints are crucial; in scenarios with sparse or irregularly distributed cameras, establishing a robust cross-view consensus becomes intrinsically challenging, which can, in turn, diminish the performance gains typically achieved by the method.
Secondly, our approach’s reliance on joint structure-view optimization, which fundamentally differs from methods that optimize only local representations (like context grids alone) or single global structures (such as basic tri-planes), introduces increased computational overhead. This is primarily due to the demands of multi-view gradient synchronization and the processing of hierarchical features, rendering the method more resource-intensive.
Consequently, addressing these aspects defines critical avenues for future work. There is a clear need for optimizing the trade-off between reconstruction fidelity and computational cost, particularly when operating under demanding capture conditions such as highly non-uniform sampling or significant viewpoint disparity. Future methodological enhancements should therefore include the development of more adaptive view sampling strategies. Additionally, creating lightweight mechanisms for enforcing cross-view consistency will be pivotal in bolstering efficiency without a concomitant sacrifice in the quality of the reconstruction.

### v4

@@ 5.5.1 Effectiveness of Cross-Structure Collaboration Module
Figure 8: Qualitative comparison of ablation study on CVPM. Red & green squares highlight the visual differences for clearer comparison. By reconstructing the fine-grained structures in two scenes, such as the steel frame and signboards in the first row of images, and lane markings and street lamps in the second row of images, we demonstrate the superiority of the CVPM and validate the effectiveness of its individual components. Please zoom in to view detailed results.
5.5.2 Effectiveness of Cross-View Pruning Mechanism
In this section, we conduct ablation studies on the CVCL and prune procedure in our CVPM. As shown in Table. 6, integrating key components enables the model to achieve optimal performance, with particularly significant improvements in SSIM metrics. Qualitative results in Fig. 8 demonstrate that incorporating the components significantly enhances rendering quality by improving both detail clarity (e.g. street lamp) and artifact suppression.
6 Conclusion and future work
In this paper, we introduce SplatCo, a structure-view collaborative framework specifically designed to address the scalability and fidelity challenges in rendering large-scale unbounded scenes. To overcome the inherent trade-off between global coherence and local details, we first propose the CSCM, which effectively harmonizes global tri-plane representations with local context grids via a novel hierarchical compensation strategy. Additionally, we implement the CVPM to mitigate Gaussian redundancy and visual artifacts, ensuring that only geometrically consistent primitives are retained. Finally, we incorporate the SVC module, which synergizes view gradients with structural gradients to robustly regularize Gaussian attributes. Extensive experiments on 13 diverse large-scale scenes demonstrate the superiority of our approach in rendering quality and detail preservation compared to state-of-the-art methods.
Future Work: In the future, we will explore reconstructing scenes with extreme structural characteristics, such as extensive textureless planes, by incorporating more robust geometric priors to enhance structural effectiveness, which is a worthwhile direction for achieving robust rendering of various challenging scenes.